OpenStudy (anonymous):

WILL GIVE MEDAL!!!!!!!!!!! Directions: Drag the tiles to the correct boxes to complete the pairs. Let A = 2, B =2/5 , and C =3/8 . Find the value of each expression below. 2 1/40 1 3/20 1 39/40 1 7/40 2 - 2/5 + 3/8 <---------> 2 + 2/5 - 3/8 <-------> 2 x 2/5 + 3/8 <--------> 2 x 3/8 + 2/5 <-------->

OpenStudy (anonymous):

@CountryGurl15

OpenStudy (anonymous):

@GreenCat

OpenStudy (anonymous):

@cupcakerain

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@michele_laino

OpenStudy (anonymous):

I know it is in Biology. I accidentally did that. :)

OpenStudy (michele_laino):

we have : 2 1/40 = 81/40

OpenStudy (anonymous):

Ok

OpenStudy (michele_laino):

now we have to search for the right computation which give us 81/40

OpenStudy (anonymous):

Ok

OpenStudy (michele_laino):

for example, I will compute the first expression:

OpenStudy (anonymous):

Ok

OpenStudy (michele_laino):

oops.. I have made an error

OpenStudy (anonymous):

Ok it is fine

OpenStudy (michele_laino):

\[2 - \frac{2}{5} + \frac{3}{8} = \frac{{2 \times 40 - 2 \times 8 + 3 \times 5}}{{40}} = \frac{{80 - 16 + 15}}{{40}} = ...?\]

OpenStudy (anonymous):

Ok it may take a minute so...

OpenStudy (anonymous):

\[\frac{ 79 }{ 40 }\]

OpenStudy (michele_laino):

ok! So the first computation doesn't match with 81/40

OpenStudy (anonymous):

Right

OpenStudy (michele_laino):

then I try to compute the second expression, as follows:

OpenStudy (anonymous):

Ok

OpenStudy (michele_laino):

\[2 + \frac{2}{5} - \frac{3}{8} = \frac{{2 \times 40 + 2 \times 8 - 3 \times 5}}{{40}} = \frac{{80 + 16 - 15}}{{40}} = ...?\]

OpenStudy (anonymous):

\[\frac{ 81 }{ 40 }\]

OpenStudy (michele_laino):

ok! we have found our first match

OpenStudy (anonymous):

YAY!!!!

OpenStudy (anonymous):

So A - B + C is 2 1/40?

OpenStudy (michele_laino):

no, A+B-C = 2 1/40

OpenStudy (anonymous):

Oh ok so lets work on the next one

OpenStudy (michele_laino):

now we have: \[1{\text{ }}3/20{\text{ }} = 1 + \frac{3}{{20}} = \frac{{23}}{{20}}\]

OpenStudy (anonymous):

Is the first one 1 39/40 for A - B + C

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

YAY!!! Ok so the last two

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

I will compute the third expression as follows: \[2 \times \frac{2}{5} + \frac{3}{8} = \frac{4}{5} + \frac{3}{8} = \frac{{4 \times 8 + 3 \times 5}}{{40}} = \frac{{32 + 15}}{{40}} = ...?\]

OpenStudy (anonymous):

\[\frac{ 47 }{ 40 }\]

OpenStudy (michele_laino):

ok! Nevertheless it is not what we are searching for

OpenStudy (anonymous):

Right so it would be the answer to which one?

OpenStudy (michele_laino):

I try to compute the fourth expression:

OpenStudy (anonymous):

Ok

OpenStudy (michele_laino):

\[2 \times \frac{3}{8} + \frac{2}{5} = \frac{6}{8} + \frac{2}{5} = \frac{{6 \times 5 + 2 \times 8}}{{40}} = \frac{{30 + 16}}{{40}} = ...?\]

OpenStudy (anonymous):

\[\frac{ 46 }{ 40}\]

OpenStudy (michele_laino):

ok! We have found our third match, since we can write: